# Self interest cumulative subtraction games

### v1

@@ 3 Unilateral tie-breaking deviation and the proof of main result
To prove the equality in item (ii), observe that, if $s^{\prime}=s^{\prime\prime}$, then the equality holds by induction, since, by item (3) in the statement, we may replace $o^{1}_{{\scriptscriptstyle{\rm FvF}}}(h-s^{\prime})$ with $o^{1}_{{\scriptscriptstyle{\rm AvF}}}(h-s^{\prime})$. We study two cases when the actions differ.
Case 1, $s^{\prime}>s^{\prime\prime}$: Suppose that $s^{\prime}=s_{1}$ and $s^{\prime\prime}=s_{2}$, corresponding to Alice’s choices, playing from $h$ in FvF and FvA, respectively, and let $H_{1}=h-s_{1}$ and $H_{2}=h-s_{2}$. By induction we may identify $o^{1}_{{\scriptscriptstyle{\rm FvF}}}(H_{1})=o^{1}_{{\scriptscriptstyle{\rm AvF}}}(H_{1})$, and study both options in the same convention, say AvF. Thus, it suffices to prove that
@@
First we prove that Bob cannot do strictly better on the smaller heap $H_{1}=H_{2}-\Delta$ in PSPE. This follows because he can mimic (couple) all moves from the smaller heap to the larger heap. If Alice at some stage plays $s_{1}$ on $H_{2}$ and $s_{2}$ on $H_{1}$, then this coupled move pair compensates exactly the initial $\Delta$ difference, so the heaps are equalized. And, given PSPE play by Alice on both heaps, since they follow the same tie-breaking convention (2) Bob’s mimic strategy obviously remains a valid strategy. (If Alice plays a smaller move on $H_{2}$, but the larger one on $H_{1}$, Bob’s strategy remains feasible.)
Now we show that Bob cannot do strictly better on the larger heap $H_{2}$. Alice sacrificed in her first PSPE move on $H_{2}$, so we know that she does not have a parity disadvantage on that heap. Let us argue first in case of dominance. By Lemma 5, Lemma 6 (ii) applies, and so the initial $\Delta$ loss will be compensated by $\Delta$ in the endgame. In case of non-dominance, Lemma 6 (i) might apply. But in this case $\Delta<s_{2}$, so Bob’s moves on the larger heap can be coupled on the smaller heap, because Alice’s share of the larger heap increased by $s_{2}$ due to parity advantage.
@@
Concerning the two last items in the statement (so $X={\rm AvF}$), it suffices to prove that
@@
Combine Lemmas 7 and 8. ∎
In case of dominance, we arrive at a stronger version of the main theorem.
@@
Consider $S=\{s_{1},s_{2}\}$, in the dominant regime when $2s_{2}\leq s_{1}$. Then there is no heap $h$ for which the $\rm FvF$} and $\rm AvA$} settings have a discrepancy. In more generality, for any heap size $h$, and any $X,Y\in\tau$, $o_{X}(h)=o_{Y}(h)$, i.e.,
@@
Consider the $H_{2}$ branch, where the previous player (Alice) has made a sacrifice. Let there be strict dominance. Thus we may assume that we are in case Lemma 6 (i), which implies that Alice lost $\Delta$ in her first move, and regained $\Delta$ in her last move. Thus, Bob’s PSPE utility remains the same as in the $H_{1}$ branch, since he does not gain anything from Alice’s sacrifice, irrespective of tie-breaking rules.
In case of the balanced regime, both Lemma 6 (i) and (ii) are possible, but in either case the argument is the same as in the previous paragraph, since there is no change in Alice’s utility in both cases. ∎
@@
During our investigation into unilateral deviation in the two-action case, we additionally investigated differences between different two-action settings (given a fixed subtraction set and initial heap size). Primarily, we investigated the discrepancies between the symmetric friendly game (where both Alice and Bob are friendly), and the symmetric antagonistic game (where both Alice and Bob are antagonistic), i.e. between the $\rm FvF$} and $\rm AvA$} games in the two-action settings. The first such instance is with $S=\{3,5\}$ and $x=14$, summarized in Table 1.
Observation 10.
Consider the $\rm FvF$} and $\rm AvA$} self-interest games for $S=\{s_{2},s_{1}\}$, where $s_{2}<s_{1}$. We tested to find if there existed a heap size $x$ where the outcomes of the two games (i.e., the difference in the players’ final cumulations) were different - i.e., a friendly/antagonistic discrepancy where $o^{1}_{{\scriptscriptstyle{\rm AvA}}}(x)-o^{2}_{{\scriptscriptstyle{\rm AvA}}}(x)\neq o^{1}_{{\scriptscriptstyle{\rm FvF}}}(x)-o^{2}_{{\scriptscriptstyle{\rm FvF}}}(x)$. Our results are summarized in Figure 1.
Figure 1: Points $(s_{2},s_{1})$ where there exists a heap with a friendly/antagonistic discrepancy for $S=\{s_{2},s_{1}\}$, $s_{2}<s_{1}$. Colors indicate similarly-sized heaps where the outcomes differ. An online version can be found here.
We first note two distinct features of this graph: (1) that there is a large region above $b=2a$ where no discrepancies were found, (2) that there are specific subtraction sets in the region below $b=2a$ where there are no discrepancies - and that these sets are integer points along lines radiating from the origin.
@@
Consider $S=\{s_{2},s_{1}\}$ where $\frac{s_{1}}{s_{2}}=\frac{k+1}{k}$ for some $k\in\mathbb{N}$. Then, there is no heap $x$ where the $\rm FvF$} and $\rm AvA$} self-interest games have a discrepancy.
Further analysis of the heaps where the two games diverge led us to conjecture the following:
@@
Consider $S=\{s_{2},s_{1}\}$ in the non-dominant regime where $s_{1}<2s_{2}$, and $\frac{s_{1}}{s_{2}}\neq\frac{k+1}{k}$ for some $k\in\mathbb{N}$. Then there exists a heap $x$ where the $\rm FvF$} and $\rm AvA$} games have a discrepancy. The first such heap can be defined by:
$x=(i+2)s_{2}+iks_{1}$
for $i\in\mathbb{N}$ satisfying $\frac{i+1}{i}>\frac{s_{1}}{s_{2}}>\frac{i+2}{k+i}$.
When we extend the methodology used in Figure 1 to $S=\{s_{3},s_{2},s_{1}\}$, where $s_{3}<s_{2}<s_{1}$, a number of new trends emerge, summarized in Figure 2.
Figure 2: Points $(s_{3},s_{2},s_{1})$ where there exists a heap with a $\rm FvF$$\rm AvA$} discrepancy for $S=\{s_{3},s_{2},s_{1}\}$, with $s_{3}<s_{2}<s_{1}$. Color indicates the value of $s_{3}$ for readability. An online rotatable version can be found here.
@@
Consider $S=\{s_{3},s_{2},s_{1}\}$ where $s_{1}\geq 2s_{2}$. Then, there does not exist a friendly/antagonistic discrepancy at any heap $x$ when there are no discrepancies for $S=\{s_{3},s_{2}\}$ for any heap $x$.
@@
Consider $S=\{s_{3},s_{2},s_{1}\}$ where $s_{1}=s_{3}+s_{2}$ - i.e., the additive scenario. Then, there exists a heap $x$ with a friendly/antagonistic discrepancy, if and only if $\frac{s_{2}}{s_{3}}\neq k$ for any natural number $k$. The first such heap is:
$x=(i+2)s_{2}+is_{1}=is_{3}+(2i+2)s_{2}$
for $i\in\mathbb{N}$ satisfying $i+1>\frac{s_{2}}{s_{3}}>i$.
@@
Let us consider zero-sum scoring play subtraction games. In this setting, Alice’s cumulations add to a common score, and Bob’s cumulations subtract from it. The outcome function in this setting is defined to be:
@@
In both zero-sum scoring play and antagonistic self-interest play, the game is built on a shared heap, so at a structural level one player’s gain reduces the other’s potential. The difference is that in zero-sum this opposition is explicit, Alice’s utility is defined as Bob’s loss. In $\rm AvA$}, by contrast, opposition arises only at tie-breaking: outside of indifference, each player maximizes their own tally without regard to the other. Thus, our exploration into discrepancies in the two-action setting naturally searched for games where the $\rm AvA$} and scoring play settings diverged (for a common subtraction set and initial heap). Consider the $\rm AvA$} self-interest game and the zero-sum scoring play game for $S=\{s_{2},s_{1}\}$, where $s_{2}<s_{1}$. However, the graph produced (using the same methodology as Figure 1) was completely empty, which we posit is a larger trend.
@@
Given any two-action subtraction set, for all heap sizes $h$, $o_{\textrm{zs}}(h)=o_{{\scriptscriptstyle{\rm AvA}}}^{1}(h)-o_{{\scriptscriptstyle{\rm AvA}}}^{2}(h)$.
Then, consider the same scenario but with $S=\{s_{3},s_{2},s_{1}\}$ where $s_{3}<s_{2}<s_{1}$. This time, discrepancies did emerge summarized in Figure 3. These discrepancies had clear planes bounding them, summarized in the conjecture below.
Figure 3: Points $(s_{3},s_{2},s_{1})$ where there exists a heap that causes a discrepancy between the $\rm AvA$} and zero-sum games for $S=\{s_{3},s_{2},s_{1}\}$. Color indicates the value of $s_{3}$ for readability. All cases where $s_{3}\not<s_{2}\not<s_{1}$ are ignored. An online rotatable version can be found here.
@@
Consider $S=\{s_{3},s_{2},s_{1}\}$ where $s_{3}<s_{2}<s_{1}$. Then, there will be no discrepancy between the zero-sum scoring play and antagonistic self-interest games if $s_{2}\leq 2s_{3}$ or if $s_{1}\geq s_{2}+s_{3}$.
We study these conjectures in a future paper [28].
@@ 6 Some more open problems
We have tested 200 random subtraction sets for each size $3\leq|S|\leq 10$, with $\max S\leq 25$. For $x\leq 300$, we have detected no positive outcome discrepancy, for either player, going from the setting FvF to AvA. This is in support of Conjecture 3.
@@
How large discrepancy ${\rm FvF}\rightarrow{\rm AvA}$ (in absolute value) is possible as a function of $\max S$?
Conjecture 17.
For any $S$, and any two tie-breaking conventions, the discrepancies are bounded.
It is known that zero-sum cumulative subtraction games are eventually periodic, with period $2\max S$ (independently of the size of $S$).
@@
Is it true that the discrepancies of self-interest subtraxtion are eventually periodic, with the period length $2\max S$?
@@
Provide a bound of the pre-period length as a function of $\max S$.
@@
Recall, that by our notation, in the first case, the starting player deviates, while in the second case the player who does not start deviates. So we have an asymmetric situation with respect to tie-breaking rules.
@@
Is it true that no player can have positive discrepancy, by going from AvF or FvA to AvA?
We have looked into max/min (greedy/sacrifice) as alternative tie-breaking convetions for indifferent actions. But player utilities, for general subtraction sets, appear mixed up, with no clear preference order between pairs of tie-breaking rules, across all positions.
Problem 7.
Is there any pair of deterministic tie-breaking rules that bound all other ones, such that PSPE utility is maximized for both players across all positions?
7 Related work
An important design choice in our framework concerns deterministic tie-breaking: how players resolve indifference between equally optimal moves. In classical combinatorial games, such ambiguity does not arise; perfect information induces unique optimal values. By contrast, our self-interest model naturally produces local indifferences, making tie-breaking an essential component of pure subgame-perfect equilibrium (PSPE) computation.
@@ 7 Related work
We resolve this by assigning each player a deterministic tie-breaking identity—friendly or antagonistic—applied uniformly throughout the game tree. One may likewise impose a deterministic max/min (greedy/sacrifice) choice. This allows backward induction to proceed unambiguously, but raises subtle philosophical and computational issues. While the legal move set is simple and local (e.g., subtract $3$ or $5$), the choice between friendly versus antagonistic (or greedy vs. sacrificial) tie-breaking can depend on the entire continuation, and hence requires full recursive evaluation. In this sense, a tie-breaking rule is both a behavioral commitment and a computational burden. In line with combinatorial game theory tradition, we assume perfect players capable of such reasoning. (In a purely recreational setting, a self-interest subtraction game should of course be played without any such commitment.)
@@
This design contrasts with many settings in economics, AI, and multi-agent systems, where tie-breaking rules are typically exogenous, randomized, or fixed to promote fairness or efficiency. In our setting, the tie-breaking rule is part of the player’s identity; it is not merely a local resolver of ambiguity, but prompts questions about efficiency in a global sense, as different deterministic conventions can be compared across positions and instances. A natural question is whether there exists a pair of deterministic tie-breaking rules that jointly outperform any other pair for both players.
@@
Across several domains, tie-breaking is often treated as a technicality, yet small local choices can propagate to global effects. In voting, different schemes (e.g., parallel-universes tiebreaking, PUT) alter strategies and collective outcomes [20]. In fair division, when multiple near-fair allocations exist, deterministic selection among them can shift welfare and perceived fairness [29]. In tournaments, stable refinements (e.g., minimal stable sets) show how resolving cycles and indifference determines which winners emerge [10]. Recent studies of simple stochastic games emphasize multiple objectives ordered lexicographically; this ordering resolves indifference and shapes equilibrium paths [12]. That said, the impact of tie-breaking is model-dependent: in well-behaved private-value auctions, equilibria can be invariant to the rule [22]. Overall, our results fit a broader pattern in which deterministic conventions for resolving indifference create system-level consequences, while also highlighting boundaries where such rules have no effect.
In computational social choice, a voting rule must specify what happens when two (or more) candidates are exactly tied. Deterministic conventions such as lexicographic priority versus random-candidate tie-breaking are often treated as “mere details”. However, recent work shows that the choice of tie-breaker has global consequences: it can change equilibria, manipulation incentives, and welfare. In particular, [5] analyze standard rules and prove that neither deterministic convention dominates across instances: each can make elections more or less manipulable and can change price-of-anarchy bounds. Thus, resolving local indifference (who wins a tie) re-routes the strategic landscape.
In sequential social-learning models (wherein agents act one-by-one, each observing predecessors’ actions and a private signal), an informational cascade arises when agents rationally ignore their signals and follow the crowd; learning may then fail even as more signals arrive [8]. A knife-edge occurs when public information and a private signal exactly balance and the agent is indifferent; [27] show that the existence and properties of cascades can hinge on this local tie-breaking convention. Relaxing the usual “if indifferent, pick action AA” assumption can admit equilibria without cascades. Again, a seemingly harmless local choice has global repercussions: whether society asymptotically learns the true state.

### v2

@@ 3 Unilateral tie-breaking deviation and a proof of the main result
To prove the equality in item (ii), observe that if $s^{\prime}=s^{\prime\prime}$, then the equality holds by induction, since, by item (3) in the statement, we may replace $o^{1}_{{\scriptscriptstyle{\rm FvF}}}(h-s^{\prime})$ with $o^{1}_{{\scriptscriptstyle{\rm AvF}}}(h-s^{\prime})$. We study two cases when the actions differ.
Case 1, $s^{\prime}>s^{\prime\prime}$: Suppose that $s^{\prime}=s_{1}$ and $s^{\prime\prime}=s_{2}$, corresponding to Alice’s choices, playing from $h$ in FvF and FvA, respectively, and let $H_{1}=h-s_{1}$ and $H_{2}=h-s_{2}$. By induction, we may identify $o^{1}_{{\scriptscriptstyle{\rm FvF}}}(H_{1})=o^{1}_{{\scriptscriptstyle{\rm AvF}}}(H_{1})$, and study both options in the same convention, say AvF. Thus, it suffices to prove that
@@
First, we prove that Bob cannot do strictly better on the smaller heap $H_{1}=H_{2}-\Delta$ in PSPE. This follows because he can mimic (couple) all moves from the smaller heap to the larger heap. If Alice at some stage plays $s_{1}$ on $H_{2}$ and $s_{2}$ on $H_{1}$, then this coupled move pair compensates exactly the initial $\Delta$ difference, so the heaps are equalized. And, given PSPE play by Alice on both heaps, since they follow the same tie-breaking convention (2), Bob’s mimic strategy obviously remains a valid strategy. (If Alice plays a smaller move on $H_{2}$, but the larger one on $H_{1}$, Bob’s strategy remains feasible.)
Now we show that Bob cannot do strictly better on the larger heap $H_{2}$. Alice sacrificed in her first PSPE move on $H_{2}$, so we know that she does not have a parity disadvantage on that heap. Let us argue first in the case of dominance. By Lemma 6, Lemma 7 (ii) applies, and so the initial $\Delta$ loss will be compensated by $\Delta$ in the endgame. In case of non-dominance, Lemma 7 (i) might apply. But in this case $\Delta<s_{2}$, so Bob’s moves on the larger heap can be coupled on the smaller heap, because Alice’s share of the larger heap increased by $s_{2}$ due to parity advantage.
@@
Regarding the last two items in the statement (so $X={\rm AvF}$), it suffices to prove that
@@
Combine Lemmas 8 and 9. ∎
In the case of dominance, we arrive at a stronger version of the main theorem.
@@
Consider $S=\{s_{2},s_{1}\}$, in the dominant regime when $2s_{2}\leq s_{1}$. Then there is no heap $h$ for which the $\rm FvF$} and $\rm AvA$} settings have a discrepancy. In more generality, for any heap size $h$, and any $X,Y\in\tau$, $o_{X}(h)=o_{Y}(h)$, i.e.,
@@
Consider the $H_{2}$ branch, where the previous player (Alice) has made a sacrifice. Let there be strict dominance. Thus, we may assume that we are in case Lemma 7 (i), which implies that Alice lost $\Delta$ in her first move, and regained $\Delta$ in her last move. Thus, Bob’s PSPE utility remains the same as in the $H_{1}$ branch, since he does not gain anything from Alice’s sacrifice, irrespective of the tie-breaking rules.
In case of the balanced regime, both Lemma 7 (i) and (ii) are possible, but in either case, the argument is the same as in the previous paragraph, since there is no change in Alice’s utility in both cases. ∎
@@
We are interested in comparing outcomes between the $\rm FvF$} and $\rm AvA$} settings, specifically when the subtraction set $S$ has exactly two actions. In this scenario, the first such discrepancy between the outcomes appears when $S=\{3,5\}$ and $h=14$, as demonstrated in Table 1.
Further computer simulations of this scenario yield a list of two-action subtraction sets $S=\{s_{2},s_{1}\}$ for which there exists a heap where the $\rm FvF$} and $\rm AvA$} outcomes differ. Our observations are summarized in Figure 1.
Figure 1: Rulesets $S=\{s_{2},s_{1}\}$ (visualized as points $(s_{2},s_{1})$, with $s_{2}<s_{1}$) for which there exists a heap with a antagonistic-friendly discrepancy. The colors indicate the value of $n$ in (5) modulo 5 for the given subtraction set (e.g., the large green area indicates $n=1$, purple is $n=2$, and so on up to $n=48$). An online version can be found here.
We note some distinct features within Figure 1:
  1. (1)
    no discrepancies occur in the region above $s_{1}=2s_{2}$,
  2. (2)
    a number of discrepancies occur with subtraction sets in the region below $s_{1}=2s_{2}$, and
  3. (3)
    the subtraction sets with no discrepancy in the region below $s_{1}=2s_{2}$, occure on half-lines radiating from the origin.
The first observation is a direct consequence of Proposition 10, while the latter two observations underpin the following conjectures.
@@
Consider $S=\{s_{2},s_{1}\}$ where $\frac{s_{1}}{s_{2}}=\frac{k+1}{k}$, for some $k\in\mathbb{N}$. Then there is no heap $h$ that distinguishes the outcomes of the two conventions $\rm FvF$} and $\rm AvA$}.
@@
Consider $S=\{s_{2},s_{1}\}$ in the non-dominant regime where $s_{1}<2s_{2}$, and supppose, for all $k\in\mathbb{N}$, $\frac{s_{1}}{s_{2}}\neq\frac{k+1}{k}$. Then there exists a heap $h$ that distinguishes the outcomes of the $\rm FvF$} and $\rm AvA$} conventions. The smallest such heap is defined by:
$h=(n+2)s_{2}+ns_{1},$ (5)
where $n\in\mathbb{N}$ satisfies $\frac{n+1}{n}>\frac{s_{1}}{s_{2}}>\frac{n+2}{n+1}$.
Note that, in Conjecture 12, the existence and uniqueness of $n$ is clear, namely $n=\left\lfloor\frac{s_{2}}{s_{1}-s_{2}}\right\rfloor$. When we extend the methodology used in Figure 1 to three-action games $S=\{s_{3},s_{2},s_{1}\}$, we use the convention $s_{3}<s_{2}<s_{1}$, and a number of new trends emerge; see Figure 2.
Figure 2: Points $(s_{3},s_{2},s_{1})$ where there exists a heap with a antagonistic-friendly discrepancy for $S=\{s_{3},s_{2},s_{1}\}$, with $s_{3}<s_{2}<s_{1}$. An online rotatable version can be found here.
@@
Consider $S=\{s_{3},s_{2},s_{1}\}$, with $s_{1}\geq 2s_{2}$, and let $S^{\prime}=\{s_{3},s_{2}\}$. If there is no discrepancy between $\rm FvF$} and $\rm AvA$} in $S^{\prime}$, then there is no discrepancy in $S$.
Let us study the so-called additive rulesets.
@@
Consider $S=\{s_{3},s_{2},s_{1}\}$ where $s_{1}=s_{3}+s_{2}$. Then, there exists a heap $h$ with a discrepancy, if and only if $\frac{s_{2}}{s_{3}}\neq k$, for any $k\in\mathbb{N}$. The smallest such heap is:
$h=(n+2)s_{2}+ns_{1}=ns_{3}+(2n+2)s_{2},$
for $n\in\mathbb{N}$ satisfying $n+1>\frac{s_{2}}{s_{3}}>n$.
@@
Let us consider zero-sum scoring play subtraction games. In this setting, Alice’s cumulations add to a common score, and Bob’s cumulations subtract from it. The outcome function [5] in this setting is:
@@
Recall that in antagonistic tie-breaking, both players minimize their opponent’s utility whenever there is an indifference. This may appear similar to the zero-sum case where the players are always working against each other. In this spirit we searched for games where the $\rm AvA$} and scoring play settings diverge, but in the case of two-action games, we did not find any discrepancies.
@@
Consider any $S=\{s_{2},s_{1}\}$. Then, for all heap sizes $h$, $o_{\textrm{zs}}(h)=o_{{\scriptscriptstyle{\rm AvA}}}^{1}(h)-o_{{\scriptscriptstyle{\rm AvA}}}^{2}(h)$.
In the case of three-action games discrepancies do emerge; see Figure 3. These discrepancies appear bounded by planes.
Figure 3: Points $(s_{3},s_{2},s_{1})$ where there is a discrepancy between $\rm AvA$} and zero-sum games for $S=\{s_{3},s_{2},s_{1}\}$, with $s_{3}<s_{2}<s_{1}$. An online rotatable version can be found here.
@@
Consider $S=\{s_{3},s_{2},s_{1}\}$ where $s_{3}<s_{2}<s_{1}$. Suppose that $s_{2}\leq 2s_{3}$ or $s_{1}\geq s_{2}+s_{3}$. Then there is no discrepancy between zero-sum scoring and antagonistic self-interest play.
@@ 6 Some more open problems
We have tested 200 random subtraction sets for each size $3\leq|S|\leq 10$, with $\max S\leq 25$. For $h\leq 300$, we have detected no positive outcome discrepancy, for either player, going from the setting FvF to AvA. This is in support of Conjecture 4.
Conjecture 17.
For any $S$ and any two tie-breaking conventions, the discrepancies are bounded.
@@
How large a discrepancy ${\rm FvF}\rightarrow{\rm AvA}$ (in absolute value) is possible as a function of $\max S$?
It is known that zero-sum cumulative subtraction games are eventually periodic, with a period of $2\max S$ (independent of the size of $S$).
@@
Is it true that the discrepancies of self-interest subtraction are eventually periodic, with the period length $2\max S$?
@@
Provide a bound on the pre-period length as a function of $\max S$.
@@
Recall that by our notation, in the first case, the starting player deviates, while in the second case, the player who does not start deviates. So we have an asymmetric situation with respect to tie-breaking rules.
@@
Is it true that no player can have a positive discrepancy by going from AvF or FvA to AvA?
7 Discussion and related work
In classical combinatorial games, such as the normal-play convention, any indifference between options implies indifference for both players; perfect information induces unique optimal outcomes. By contrast, our self-interest model can produce local indifferences with discrepancies in the result for the other player. And thus, the choice of tie-breaking mechanism becomes essential in any pure subgame-perfect equilibrium (PSPE) computation. Our interest in staying close to combinatorial game theory lies in studying the deterministic backwards-induction algorithms induced by specific choices of deterministic tie-breaking mechanisms.
@@ 7 Discussion and related work
Our approach has been to assign each player a deterministic tie-breaking identity, friendly or antagonistic, applied uniformly throughout the game tree. One may likewise impose a deterministic max/min (greedy/sacrifice) choice. For example, if $S=\{3,5\}$, then a greedy deterministic tie-breaking rule would be to always choose $5$ in case of indifference. This would allow backward induction to proceed unambiguously. Such a tie-breaking rule may seem simpler; however, the question of whether there is indifference between the options still requires a complete backward induction computation. The choice between friendly versus antagonistic (or greedy vs. sacrificial) tie-breaking depends on the entire continuation and hence requires full recursive evaluation. In this sense, either tie-breaking rule is both a behavioral commitment and a computational burden. In line with the tradition of combinatorial game theory, we assume perfect players capable of such reasoning. In a purely recreational setting, any self-interest subtraction game should, of course, be played without any such commitment.
This perspective contrasts with many settings in economics, AI, and multi-agent systems, where tie-breaking rules are imposed exogenously, sometimes randomized to ensure fairness, or fixed for predictability or efficiency.
For instance, in a standard auction or matching model, if two agents submit identical bids or preferences, the model typically resolves the tie by randomization or by a fixed priority order specified in advance; the agents themselves do not control this rule, and it is not typically intended to affect the strategic structure beyond selecting one outcome among otherwise equivalent ones. Similarly, in many planning or learning algorithms, when multiple actions have equal value estimates, a fixed lexicographic rule is used to select one action, ensuring the algorithm is well-defined, rather than modelling any strategic preference of the agent. A related illustration arises in school choice: when applicants have the same score for a seat, a deterministic priority list systematically favors the same students, whereas random tie-breaking treats equally qualified applicants symmetrically.
@@
In our setting, the tie-breaking rule is part of the player’s identity; it is not merely a local resolver of ambiguity, but prompts questions about comparative efficiency in a global sense, as different deterministic conventions can be compared across positions and instances. A natural question is whether there exists a pair of deterministic tie-breaking rules that jointly outperform any other pair for both players. We have tested both max/min and friendly/antagonistic, and we have not found any such pair.
@@
Across several domains, tie-breaking is often treated as a technicality, yet small local choices can propagate to global effects. In each of the following examples, a locally innocuous rule for resolving indifference alters the structure or outcomes of the model as a whole. In voting, different schemes (e.g., parallel-universes tiebreaking, PUT) alter strategies and collective outcomes [6]. In fair division, when multiple near-fair allocations exist, deterministic selection among them can shift welfare and perceived fairness [13]. In tournaments, cycles can make winner selection underdetermined, and deterministic refinements such as minimal stable sets select different winner sets on different instances, so the choice of convention is not neutral [4]. That said, the impact of tie-breaking is model-dependent: in well-behaved private-value auctions, equilibria can be invariant to the rule [7]. Overall, our results fit a broader pattern in which deterministic conventions for resolving indifference have system-level consequences, while also highlighting boundaries where such rules have no effect.